OpenStudy (anonymous):

Big test coming up! Need help with 16 Geometry Problems! :) Please help. ~In Dire Need of Help

OpenStudy (jack1):

geometry...that's all about rocks and volcanos an stuff...yeah?

OpenStudy (anonymous):

First seven questions...

OpenStudy (anonymous):

Geometry is the study of shapes

OpenStudy (jack1):

jks, sorry ;P

OpenStudy (jack1):

Q1: volume = area of base x ht

OpenStudy (jack1):

shape 1 base = rectangle formula for area of rectangle =...? @djibben615 ?

OpenStudy (anonymous):

yep. I got question number 1. I need help with rest :)

OpenStudy (anonymous):

@mathmale @amistre64

OpenStudy (mathmale):

Go right ahead. Make your own choice of answer and explain your reasoning.

OpenStudy (jack1):

Q5 - volume of sphere = \[ \large \frac 43 \pi r^3\] and \[\large r = d/2\] so how does change in d affect change in V ?

OpenStudy (jack1):

hint... 18:24 = 3:4 ratio in diameters

OpenStudy (anonymous):

ah. ok. do I do R=d/2 twice for both snowballs?

OpenStudy (anonymous):

So it is 8 times bigger?????

OpenStudy (jack1):

q4: vol of cylinfer = area of base x ht = pi r ^2 x ht cyl 1 = pi r ^2 x ht = pi 1^2 x 10 = 10pi inches sqrd or volume cyl 2 10pi= pi r ^2 x ht 10pi= pi 2^2 x ht 10 = 4x ht ht = 10/4

OpenStudy (jack1):

q3... i dont know jack about imperial measurements... whats an ounce?

OpenStudy (jack1):

Q3 volume of barrel = area of base x ht = pi r ^2 x ht = pi 4.25^2 x 11 = 624.195 cubic inches

OpenStudy (jack1):

is 8 ounces a gallon?

OpenStudy (anonymous):

um...no its not

OpenStudy (jack1):

q6: 2 long, 4 wide, 6 high vol of single box = xyz vol of shrink wrap = 2x*4y*6z = xyz*48 so 48 times the volume

OpenStudy (jack1):

what's the relationship of gallons to OZ then, as we use litres down here?

OpenStudy (jack1):

q7... no idea what that is, sorry

OpenStudy (jack1):

back to q 1... u know the rule for area of a hexagon...?

OpenStudy (anonymous):

128 ounces in one gallon

OpenStudy (jack1):

ahh, cool... how many cubic inches in a gallon?

OpenStudy (anonymous):

not sure.

OpenStudy (anonymous):

Could we start at question 2? I still never got the answer completely. How many times bigger was it again?

OpenStudy (jack1):

apparently its 231... so lets go with that \[\large 624.195in^3 \times \frac {1 gal}{231in^3} \times \frac{128 oz}{1 gal} \times \frac {1 cup}{8oz} = ?\]

OpenStudy (jack1):

\[\large =\frac{79896.96}{1848} = 43.24 cups\]

OpenStudy (jack1):

Q2.. sure

OpenStudy (jack1):

vol of snowball 1 vol of sphere = 4/3 pi r ^3 r = diameter/2 = 18/2 = 9 so vol of sphere = 4/3 pi r ^3 vol of sphere = 4/3 pi 9 ^3 vol of sphere = 3053.63 (or 972pi) vol of snowball 2 vol of sphere = 4/3 pi r ^3 r = diameter/2 = 24/2 = 12 so vol of sphere = 4/3 pi r ^3 vol of sphere = 4/3 pi 12 ^3 vol of sphere = 7238.23 (or 2304pi) ratio volume 1 to volume 2 = 972 : 2304 = 27 to 64 ratio radius 1 to radius 2 = 9 : 12 so \[\large \frac 9{12} \times x = \frac {27}{64}\] therefore x = 2/3

OpenStudy (jack1):

but that's only comparing radius to volume \[\large \frac 9{12} \times x = \frac {27}{64}\] \[\large \frac {d1}{d2} \times x = \frac {27}{64}\] \[\large \frac {18}{24} \times x = \frac {27}{64}\] \[\large x = \frac {9}{16}\]

OpenStudy (anonymous):

So....the answer is 0.6?

OpenStudy (jack1):

so 18/24 = 3/4 we started with a 3 : 4 ratio of diameter of snowball 1 to diameter of snowball 2 now we multiply by 9 to 16 ratio to get the volume 9 to 16 ratio = (3/4)^2 so 3/4 x 3/4 x 3/4 = 27/64 so relationship between volume and diameter is ratio of diameters...cubed (to power of 3)

OpenStudy (anonymous):

So its 0.5?

OpenStudy (jack1):

think about it this way: first snowball = 3cm across 2nd snowball = 15 cm across ratio of length 1 to length 2 = 3: 15 = 1 : 5... new snowball is 5 times the length of old therefore if ratio of new to old is 5: 1 for diameters, ratio of volumes should be 5^3 to 1 so 5x5x5 = 125 x times volume

OpenStudy (jack1):

proof: vol of snowball 1 = 36pi vol of snowball 2 = 4500pi 36 : 4500 = 1 : 125 = 125 times the size = 5^3 times the size

OpenStudy (jack1):

no, not .6 or .5 find ratio of lengths cube that ratio = ratio of volumes

OpenStudy (anonymous):

it says to round to the nearest tenth...

OpenStudy (jack1):

ok, then ratio = 4 to 3 (4/3 cubed) = 64/27 = 2.37 times more volume = 2.4 (rounded)

OpenStudy (jack1):

so arnies snowball is 2.4 times the volume of robertos snowball

OpenStudy (anonymous):

AH gotcha!

OpenStudy (jack1):

sweeeet :)

OpenStudy (jack1):

so type back yes or no if u got them down: 1a?

OpenStudy (anonymous):

can we do number #2?

OpenStudy (anonymous):

So I have numbers 1, 3, 4, & 5 done. If we finish number 2, we can move on to the last 8 problems.

OpenStudy (jack1):

area of a hexagon = (3 x sqrt 3 )/ 2 times side sqrd

OpenStudy (jack1):

\[\Large Area = \frac {3 \times \sqrt 3}{ 2} \times s^2\]

OpenStudy (jack1):

volume = area of base x ht u can calculate from here?

OpenStudy (anonymous):

um....I may need some assistance. What does the s stand for again in your equation?

OpenStudy (jack1):

side, so the side length of one of the sides of the hexagon

OpenStudy (anonymous):

ah sorry. I really am not this dumb...just struggle with this chapter.

OpenStudy (jack1):



OpenStudy (jack1):

all good, dumb ppl dont ask questions, so ur a step ahead of the rest in my view

OpenStudy (anonymous):

thanks! :) So, how do I find the side length? I can calculate the equation but not sure how to find the side length.

OpenStudy (jack1):

side length should be given... in this case it's 8

OpenStudy (anonymous):

oH! I feel really stupid. I didn't notice the base was the same length. :)

OpenStudy (jack1):

hall good, chill hey ;)

OpenStudy (anonymous):

\[Area = 3 \times \sqrt{3} \div 2 \times 8^{2}\] Is this correct? And this will give me my answer?

OpenStudy (jack1):

that will only give u the areaa of the base... to get volume u need to times it by how long/heigh it it

OpenStudy (anonymous):

So 166.28? That's my answer for the equation. Now what is the next equation?

OpenStudy (anonymous):

I just multiply 166.28 x 8 x 20??

OpenStudy (jack1):